# Exploring Holomorphic Retracts

### v1

@@ 2.5. Fixed point sets and Retracts
Thus in particular, by the above theorem 2.1, all retracts of the domain of our interest here, namely $D:=B_{1}\times B_{2}$, are submanifolds thereof and moreover are all contractible since $B_{1}\times B_{2}$ is; as already indicated, retracts of our convex domain $D$ are also strong holomorphic deformation retracts – indeed, the linear homotopy between the identity map of $D$ and any retraction gives a homotopy in which the retract remains pointwise fixed  [^2] In general, while every (continuous) retract of a contractible topological space $X$ is a deformation retract much as any self-map of $X$ is homotopic to the identity, not every retract is a strong deformation retract. Infact, there are contractible spaces which do not admit a strong deformation retraction to any of its points, though obviously they admit a deformation retraction to each of its points. during the homotopy and for each point of time in the homotopy, the mapping is holomorphic. While this means that all retracts of $D$ are topologically trivial thereby placing strong restrictions on which complex submanifolds of $D$ can happen to be a retract, we shall see on the other hand that there is some variety as well: for instance if the factors $B_{1},B_{2}$ are either a polydisc or a Euclidean ball, there are retracts for every possible dimension which thereby have got to be holomorphically distinct. Furthermore again, not every contractible complex submanifold of $D$ passing through the origin can qualify to be a retract, for as we shall show, retracts through the origin of $B_{1}\times B_{2}$ are submanifolds in normal form in the sense that they are all graphs of holomorphic maps; as is very well-known, graphs of holomorphic functions are submanifolds which are simultaneously in parametric form as well as in level set form and the implicit function theorem allows us to put every submanifold locally in such a normal form (as graphs). In view of the foregoing theorem (and the implicit function theorem) then, retracts are obviously graphs of holomorphic maps (over some set of axes) locally; however in the specific case of retracts of $D=B_{1}\times B_{2}$ passing through the origin, we shall show that they are actually so globally i.e., each retract of the product domain $D$ is the graph of a single holomorphic map over some linear submanifold of $D$. As we shall encounter retracts in ‘graphical form’ in several cases in the sequel, we make a general definition for clarity.
@@
We can now progress towards our aforementioned general result on fully attaining $Z$ as the image of $L$ under $\rho$, the proof of which requires the subsequent lemmas which in turn follow fairly directly from the ideas in the pair of remarks 2.6 and 2.7, allowing us to skip details here.
@@
If we relax the convexity condition on $D$ to pseudoconvexity, it is still true that the Minkowski functional of $D$ and the infinitesimal Kobayashi metric of $D$ agree at the origin (cf. proposition 3.5.3 of [Jrncki_invrnt_dst] wherein counterexamples to show the failure of the equality when pseudoconvexity is dropped are also given), thereby implying that the Kobayashi indicatrix of $D$ at the origin is $D$ itself; likewise the Kobayashi indicatrix of $D_{L}$ now also a bounded balanced pseudoconvex domain in $L$ is also (a copy of) $D_{L}$ itself. The general property of contractivity of the Kobayashi metric under holomorphic mappings when applied to $\rho$ viewed as a map from $D$ to $D_{L}$, then yields that $D\rho(0)(D)\subset D_{L}$. But then as $D\rho(0)$ fixes $L\supset D_{L}$ pointwise, it follows that $D\rho(0)(D)=D_{L}$, showing that the above lemma holds in this more general case which we record separately for some convenience.
@@
Let $D$ be any bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$, $Z$ a retract passing through the origin with $\rho:D\to Z$, a retraction map. Let $L$ denote the tangent space to $Z$ at the origin and $D_{L}:=D\cap L$. Then, the linear map $D\rho(0)$ is a projection which retracts $D$ onto $D_{L}$.
In case $Z$ is already a linear subspace of $D$ to begin with, the above lemma still applies to show that a linear subspace of pseudoconvex bounded balanced domain $D$, which is apriori only known to be attainable as the image of a holomorphic (i.e., possibly non-linear) retraction map on $D$, can indeed be obtained as the image of a linear retraction map as well i.e., it is valid to say that $D_{L}$ as in the above lemma is a linear retract of $D$ without any further qualifications. We are now fully prepared to prove the highlighted general result about the equality $\rho(D_{L})=Z$ whose statement is formulated with slightly different notations than the one that we had adopted for our specific domain $D$. This is indeed the content of theorem 1.10, whose statement we lay down again here for convenience, and then prove it inspired by an idea in [Abate_Isometry].
@@
Let $D$ be any bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$, $Z$ a retract passing through the origin with $\rho:D\to Z$, a retraction map. Let $L$ denote the tangent space to $Z$ at the origin. Then, $\rho_{|D_{L}}$ is a biholomorphism mapping $D_{L}=D\cap L$ onto $Z$ and upto a linear change of coordinates, $Z$ is the graph of a (single) holomorphic function over $D_{L}$.
@@
Under the hypotheses that the boundary $\partial B$ is holomorphically extreme, it was shown in [JJ] that $Z$ is actually linear. While this hypothesis is the best possible for obtaining such a strong conclusion, it is restrictive as well for, there are plenty of balanced pseudoconvex domains whose boundaries fail to satisfy such a strong extremality condition. To give just one example whose boundary does not contain any flat pieces, which indeed satisfies the weaker $\mathbb{C}$-extremality but not the holomorphic extermality condition, consider the bounded balanced pseudoconvex domain obtained by intersecting the domain $H=\{(z,w)\in\mathbb{C}^{2}\;:\;|zw|<1/2\}$ with the standard Euclidean ball $\mathbb{B}$ in $\mathbb{C}^{2}$. This indicates how we may construct plenty of such examples.
@@
Being a graph, $Z$ is easily seen to be biholomorphic to $D_{L}$ as complex manifolds, in the setting of the above theorem. Thereby, it immediately follows that any pair of retracts through the origin in $D$, of the same dimension are diffeomorphic. Indeed, being star-shaped domains in $L$, they are diffeomorphic to $L$ – no pseudoconvexity is required here; the general fact being used here is that any star-shaped domain in $\mathbb{R}^{N}$ is $C^{\infty}$-diffeomorphic to $\mathbb{R}^{N}$ (cf. exercise $7$ of chapter $8$ of [Brocker-Janich-difftop-bk]). It is therefore of interest then (as is in Several Complex Variables), to note that such a pair of retracts (through the origin of the same balanced domain $D$ and) of the same dimension despite being mutually diffeomorphic need not be biholomorphic (indeed, they are rarely so). To see this first by what is perhaps the simplest example, consider $D=\Delta^{2}\times\mathbb{B}^{2}$; let $Z_{1}$ denote the projection of $D$ onto the span of the first pair of standard basis vectors of $\mathbb{C}^{4}$ and $Z_{2}$, the projection of $D$ onto the span of the remaining last pair of standard basis vectors of $\mathbb{C}^{4}$. Noting that $Z_{1}\simeq\Delta^{2}$, $Z_{2}\simeq\mathbb{B}^{2}$ and recalling that the ball and polydisc are biholomorphically inequivalent, we deduce the same about $Z_{1},Z_{2}$. To give another example, again as simple as possible but this time of a ball $D$ which is not biholomorphic to a product domain, in-fact with boundary both smooth as well as holomorphically extreme and, which contains pairs of linear retracts (thereby through the origin), of the same dimension that are biholomorphically inequivalent, consider
@@
where the exponents $p_{j}$’s are taken to be bigger than one (we disallow $N=2$ for that case is exceptional owing to the fact that all non-trivial retracts of convex domains in this case are at-most of dimension one and any pair of one-dimensional retracts are biholomorphic to $\Delta$). Then, for every $1<k<N$, there exist a pair of linear retracts of $\mathcal{E}$ of the same dimension $k$ which are diffeomorphic but biholomorphically inequivalent. This is easily seen for instance, by projecting $\mathcal{E}$ onto the span of different subsets of $k$-many of the standard basis vectors of $\mathbb{C}^{N}$ following the foregoing concrete examples.
@@
While the foregoing theorem 3.7, already realizes $Z=\rho(D_{L})$ as the graph of a holomorphic function over $L$, it must be noted that this does not mean that the graph of any holomorphic function over any given linear retract of $D$, is a retract: for instance, even in the simple case of $\mathbb{B}$, the standard unit ball with respect to the usual $l^{2}$-norm on $\mathbb{C}^{N}$, every retract through the origin is just a linear subspace and in fact, all retracts of $\mathbb{B}$ are affine-linear, as was first established by Rudin who then exposited it in chapter $8$ of his book [Rud_book_Fn_theory_unitball] as well. The proof of this, atleast as in [Rud_book_Fn_theory_unitball], uses the special feature of the $l^{2}$-norm that it comes from an inner product; however, this fact about the linearity of retracts through the origin is actually true for every bounded balanced convex domain with $\mathbb{C}$-extremal boundary. This is of-course the result of Vesentini which we extend and give a more elementary proof as already mentioned in the introduction i.e., we now lay down the proof of theorem 1.3.
Proof of theorem 1.3: Let $\rho:D\longrightarrow D$ be a retraction and $\rho(0)=0$. By lemma 3.6, $D\rho|_{0}$ is a projection of norm $1$ from $D$ onto $D_{L}=D\cap L$, where $L=T_{0}Z$. Note that $(D\rho|_{0})_{|_{D_{L}}}:D_{L}\longrightarrow D_{L}$ is the identity map on $D_{L}$ and that $D_{L}$ is a bounded balanced convex domain in $\mathbb{C}^{N}$. It follows from the proof of Krein-Milman theorem (the reader is referred to the book [Meggn] for the standard proof we have in mind here), that $D_{L}$ must have atleast one boundary point $\mathbb{R}$-extremal point, say $z_{0}$. As $D\rho(0)$ acts as the identity map on $D_{L}$ (thereby in particular $D\rho(0)(z_{0})=z_{0}$), it follows that there is atleast one $\mathbb{R}$-extremal point in $D\rho(0)(D_{L})$ namely $z_{0}$ itself! Theorem 3.7 part-(iii) of Mazet’s article [Mazet_Rextrm] applies to the map $\rho_{|_{D_{L}}}$ finishes the proof that $Z$ is linear.
@@
$L=\Delta^{N}\cap Y=\{(F_{1}(z),\ldots,F_{N}(z)):z\in M\}$
Observe that the $F_{j}$’s are holomorphic functions from $\Delta^{N}$ to $\Delta$ and $F_{j}(z)=G_{j}(z_{j_{1}},\ldots,z_{j_{r}})$ for some holomorphic function $G_{j}$, which is bounded by $1$. Without loss of generality assume that each $F_{i}$’s are linear, since $F(z)=(F_{1}(z),\ldots,F_{N}(z))$ restricted to $L$ is linear. So we get $F_{i}(z)=z_{i}$, if $i\in J$ and $F_{i}(z)=c_{i,j_{1}}z_{j_{1}}+\ldots+c_{i,j_{r}}z_{j_{r}}$, if $i\notin J$ for some scalars $c_{i,j},~{}j\in J$.
@@
$w_{k}:=\begin{cases}\frac{|c_{m,j_{r}}|}{c_{m,j_{r}}},\ \text{if}\ k=j_{r}~{}\text{for some}~{}r\ \text{and}\ c_{m,j_{r}}\neq 0\\ 0,\quad\quad\text{otherwise}\end{cases}$
@@
Hence we get $\sum_{j\in J}|c_{m,j}|\leq 1$.
Note that $M$ is spanned by $\{e_{j_{1}},\ldots,e_{j_{r}}\}$. Since $F$ is determined by the variables $z_{j_{1}},\ldots,z_{j_{r}}$, we get $Y$ is spanned by $\{F(e_{j_{1}}),\ldots,F(e_{j_{r}})\}$. By definition of $F_{i}$, we get
$F_{i}(e_{j_{k}})=\begin{cases}1,\quad\qquad\text{if}~{}i\in J\\ c_{i,j_{k}},\qquad\text{if}~{}i\notin J,\\ \end{cases}$
By definition of $F$, we will get
$F(e_{j_{k}})=(F_{1}(e_{j_{k}}),\ldots,F_{N}(e_{j_{k}}))=e_{j_{k}}+\sum_{i\notin J}c_{i,j_{k}}e_{i}$

### v6

@@ 2.5. Fixed point sets and Retracts
Thus in particular, by theorem 2.1 all retracts of the domain of our interest here, namely $D:=B_{1}\times B_{2}$, are submanifolds thereof and moreover are all contractible since $B_{1}\times B_{2}$ is; as already indicated, retracts of our convex domain $D$ are also strong holomorphic deformation retracts – indeed, the linear homotopy between the identity map of $D$ and any retraction gives a homotopy in which the retract remains pointwise fixed  [^3] In general, while every (continuous) retract of a contractible topological space $X$ is a deformation retract much as any self-map of $X$ is homotopic to the identity, not every retract is a strong deformation retract. Infact, there are contractible spaces which do not admit a strong deformation retraction to any of its points, though obviously they admit a deformation retraction to each of its points. during the homotopy and for each point of time in the homotopy, the mapping is holomorphic. While this means that all retracts of $D$ are topologically trivial thereby placing strong restrictions on which complex submanifolds of $D$ can happen to be a retract, we shall see on the other hand that there is some variety as well: for instance if the factors $B_{1},B_{2}$ are either a polydisc or a Euclidean ball, there are retracts for every possible dimension which thereby have got to be holomorphically distinct. Furthermore again, not every contractible complex submanifold of $D$ passing through the origin can qualify to be a retract, for as we shall show, retracts through the origin of $B_{1}\times B_{2}$ are submanifolds in normal form in the sense that they are all graphs of holomorphic maps; as is very well-known, graphs of holomorphic functions are submanifolds which are simultaneously in parametric form as well as in level set form and the implicit function theorem allows us to put every submanifold locally in such a normal form (as graphs). In view of the foregoing theorem (and the implicit function theorem) then, retracts are obviously graphs of holomorphic maps (over some set of axes) locally; however in the specific case of retracts of $D=B_{1}\times B_{2}$ passing through the origin, we shall show that they are actually so globally i.e., each retract of the product domain $D$ is the graph of a single holomorphic map over some linear submanifold of $D$. As we shall encounter retracts in ‘graphical form’ in several cases in the sequel, we make a general definition for clarity.
@@
The following theorem is the generalization of above Mazet theorem, which is also useful in this sequel.
Theorem 3.5.
(Theorem 3.7 of [Mazet_Rextrm]) Let $E$ and $F$ be Banach spaces and $B_{E},B_{F}$ be the corresponding unit ball. Consider a holomorphic map $f:B_{E}\to B_{F}$ be a holomorphic map. Then we have:
  1. i)
    If there exists $u$ of norm 1 in $E$ such that $f^{\prime}(0)\cdot u$ be $\mathbb{R}$-extremal in $\overline{B_{F}}$ then $f(0)=0$.
  2. ii)
    If, for any $u$ of norm 1 in $E,f^{\prime}(0)\cdot u$ is $\mathbb{C}$-extremal in $B_{F}$ then $f$ is a polynomial of degree at most 2.
  3. iii)
    If the two previous hypotheses are simultaneously verified, then $f$ is a linear map, $f(u)=$ $f^{\prime}(0)\cdot u$
@@
We can now progress towards our aforementioned general result on fully attaining $Z$ as the image of $L$ under $\rho$, the proof of which requires the subsequent lemmas which in turn follow fairly directly from the ideas in the pair of remarks 2.7 and 2.8, allowing us to skip details here.
@@
If we relax the convexity condition on $D$ to pseudoconvexity, it is still true that the Minkowski functional of $D$ and the infinitesimal Kobayashi metric of $D$ agree at the origin (proposition 3.5.3 of [Jrncki_invrnt_dst] wherein counterexamples to show the failure of the equality when pseudoconvexity is dropped are also given), thereby implying that the Kobayashi indicatrix of $D$ at the origin is $D$ itself; likewise the Kobayashi indicatrix of $D_{L}$ now also a balanced pseudoconvex domain in $L$ is also (a copy of) $D_{L}$ itself. The general property of contractivity of the Kobayashi metric under holomorphic mappings when applied to $\rho$ viewed as a map from $D$ to $D_{L}$, then yields that $D\rho(0)(D)\subset D_{L}$. But then as $D\rho(0)$ fixes $L\supset D_{L}$ pointwise, it follows that $D\rho(0)(D)=D_{L}$, showing that the above lemma holds in this more general case which we record separately for some convenience.
@@
Let $D$ be any balanced pseudoconvex domain in $\mathbb{C}^{N}$, $Z$ a retract passing through the origin with $\rho:D\to Z$, a retraction map. Let $L$ denote the tangent space to $Z$ at the origin and $D_{L}:=D\cap L$. Then, the linear map $D\rho(0)$ is a projection which retracts $D$ onto $D_{L}$.
As a corollary to this lemma, we not only get an alternative proof of corollary 2.33 – which was only for one-dimensional retracts – but also an improvement wherein there is no restriction on the dimension of the retract.
Corollary 3.8.
Let $D$ be a balanced pseudoconvex domain in $\mathbb{C}^{N}$. If there exists a point $p\in\partial D$ such that $D$ is not convex at $p$, then there does not exist a retract of $D$ passing through the origin in the direction of $p$.
Proof.
Assume to obtain a contradiction that there exists a retract $Z$ of $D$ passing through the origin in the direction of $p$. Let $\rho$ be any one of the retraction maps on $D$ such that ${\rm image}(\rho)=Z$. By the previous lemma 3.7, the linear projection map $\pi:=D\rho|_{0}\;:\;\mathbb{C}^{N}\to\mathbb{C}^{N}$ maps $D$ onto $D_{L}$, where $L=T_{0}Z$. Consider the affine subspace $K:=p+Ker(\pi)$ of $\mathbb{C}^{N}$. Since $D$ is not convex at $p$, there exists a point $q\in K$ such that $q\in D\cap K$. But $\pi(q)=p$, which contradicts the fact that $\pi$ maps $D$ into $D$. ∎
We return to the preceding lemma for a little clarification that saves some confusion once and for all. In case $Z$ is already a linear subspace of $D$ to begin with, the above lemma 3.7, still applies to show that a linear subspace of pseudoconvex bounded balanced domain $D$, which is apriori only known to be attainable as the image of a holomorphic (i.e., possibly non-linear) retraction map on $D$, can indeed be obtained as the image of a linear retraction map as well i.e., it is valid to say that $D_{L}$ as in the above lemma is a linear retract of $D$ without any further qualifications. We are now fully prepared to prove the highlighted general result about the equality $\rho(D_{L})=Z$ whose statement is formulated with slightly different notations than the one that we had adopted for our specific domain $D$. This is indeed the content of theorem 1.13, whose statement we lay down again here for convenience, and then prove it inspired and based on an idea in [Abate_Isometry].
@@
Let $D$ be any bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$, $Z$ a retract passing through the origin with $\rho:D\to Z$, a retraction map. Let $L$ denote the tangent space to $Z$ at the origin. Then, $\rho_{|D_{L}}$ is a biholomorphism mapping $D_{L}=D\cap L$ onto $Z$ and upto a linear change of coordinates, $Z$ is the graph of a holomorphic map over $D_{L}$ taking values in $M:=\ker\left(D\rho(0)\right)$.
@@
While the above theorem allows retracts to be non-linear and we shall indeed furnish several examples of non-product (balanced, pseudoconvex) domains possessing non-linear retracts through origin in the form of graphs, we recall that such a phenomenon is precluded by Vesentini’s theorem in the convex case. Extremality boundary assumptions as in Vesentini’s theorem are indeed necessary as we shall see while establishing a converse to his theorem later. However, the hypothesis that the entire boundary be holomorphically extreme as in the extension of Vesetini’s theorem given by proposition 1.4 in [JJ] is inconvenient for discussing interesting intermediate examples in the sequel. This is addressed by proposition 1.4 which we now establish. Though this is in essence an imitation of the proof of the aforementioned proposition in [JJ], we prefer to include some details as this is a result about retracts that we shall need later in this article wherein the aforementioned relaxation will actually be used. And, also because it provides an answer towards the natural query as to whether an extension of Vesentini’s theorem holds if we add only an extremality hypothesis just as much as is needed about the boundary, but none about convexity in the setting of the above theorem 3.9.
Proof of proposition 1.4: Choose any holomorphic retraction map $\rho:D\to D$ with $\rho(D)=Z$. We may observe by the above theorem 3.9, that the restriction $\rho|_{D_{L}}$ is a holomorphic map on $D_{L}$ whose image is precisely equal to $Z$; but we shall not need this for the rest of this proof here. Recall that $Z$ is a connected complex submanifold and the tangent space to $Z$ at the origin, is denoted by $L$; so, $L=T_{0}Z=\{v\in\mathbb{C}^{N}:\rho^{\prime}|_{D_{L}}(0)v=v\}$. Let $v\in D_{L},v\neq 0$. By our assumption on $\partial D_{L}$, it follows that the mapping
$\phi:\Delta\ni\lambda\longmapsto\lambda v/h_{D_{L}}(v)\in D_{L}.$
is the unique (modulo $Aut(\Delta)$) $K_{D_{L}}$-geodesic for $(0,v/h_{D_{L}}(v))$; here of-course $h_{D_{L}}$ denotes the Minkowski functional of the balanced domain $D_{L}\subset L$. Since $D_{L}$ is a linear subspace of $D$, $h_{D_{L}}(v)=h_{D}(v)$. So $\phi$ is the unique (modulo $Aut(\Delta)$) $K_{D}$-geodesic for the datum $(0,v/h_{D_{L}}(v))$. Note that
$K_{D}\left(\rho|_{D_{L}}\circ\phi(0);(\rho|_{D_{L}}\circ\phi)^{\prime}(0)\right)=K_{D}\left(0,v/h_{D_{L}}(v)\right)=K_{D}\left(\phi(0);\phi^{\prime}(0)\right).$
Hence $\rho|_{D_{L}}\circ\phi$ is a $K_{D}$-geodesic for $(0,v/h_{D_{L}}(v))$. Because of our hypothesis, we have that $v/h_{D_{L}}(v)$ is a holomorphic extreme point of $D$. By uniqueness of complex geodesics in $D$ (exposited as proposition 11.3.5 in [Jrncki_invrnt_dst]), it follows that $\rho|_{D_{L}}\circ\phi=\phi\circ\psi$, where $\psi\in{\rm Aut}(\Delta)$. Substituting $\psi^{-1}(h_{D_{L}}(v))$ which is a point in $\Delta$, into the above equation, we see that $v$ lies in the image of $\rho|_{D_{L}}$. Hence $D_{L}\subset Z$. Now $T_{0}Z$ and $Z$ have the same dimension. Since $Z$ is connected, it follows from the principle of analytic continuation that $Z=D_{L}$. Hence $Z$ is linear. ∎
Some remarks about the foregoing pair of results are in order.
@@
To indicate just a few concrete examples to illustrate where proposition 1.4 may be applied (but not the other results in this subject directly at-least), we first deal with domains where the holomorphic extremality condition is neither violated maximally nor minimally; examples for ‘maximal’ violation is also considered but in section 8. Among the simplest of domains satisfying the hypotheses of proposition 1.4 in this intermediate fashion, we consider the domain $\Omega_{1}:=\Delta^{2}\cap D_{q}$, where $D_{q}:=\{(z,w)\in\mathbb{C}^{2}:|z|^{q}+|w|^{q}<2^{q}\}$ for $10<q<$ 1 is a dilated copy of the $\ell_{q}$-ball. Note that $\Omega_{1}$ is a bounded balanced pseudoconvex domain and $\partial\Omega_{1}$ fails to be holomorphically extreme (in its entirety); though of-course, there exists an open piece of $\partial\Omega_{1}$, which is holomorphically extreme. A combined application of theorem 1.10, theorem 1.13 and corollary 2.33, proves that every retract passing through origin is either a linear subspace of $\Omega_{1}$ or a graph of a non-linear holomorphic function over some linear subspace of $\Omega_{1}$.
The above example was non-convex; so, next we consider a convex domain just by taking its limiting case i.e., letting $q\to 1$ in the above namely, $\Omega_{2}:=\Delta^{2}\cap D_{1}$, where $D_{1}:=\{(z,w)\in\mathbb{C}^{2}:|z|+|w|<3/2\}$ is some concrete dilated copy of the $\ell_{1}$-ball. Note that $\partial\Omega_{2}$ is not strictly convex, indeed not holomorphically extreme. Though no little piece of the boundary is $\mathbb{R}$-extremal, there exists an open piece of the boundary which satisfies the $\mathbb{C}$-extremality condition. By proposition 1.4 whose applications we are after, we obtain that every retract through the origin is either a linear subspace of $\Omega_{2}$ or a graph of a non-linear holomorphic function over some linear subspace of $\Omega_{2}$.
@@
While it is clear that without any extremality assumptions, linearity of retracts is unreasonable to be expected (polydisc being the simplest ‘counterexample’) and, holomorphic extremality seems the best possible condition for obtaining such a strong conclusion as in the proposition 1.4 just discussed, it is restrictive as well. For, there are plenty of balanced pseudoconvex domains whose boundaries fail to satisfy such a strong extremality condition. To give just one example whose boundary does not contain any flat pieces, which indeed satisfies the weaker $\mathbb{C}$-extremality but not the holomorphic extermality condition, consider the bounded balanced pseudoconvex domain obtained by intersecting the domain $H=\{(z,w)\in\mathbb{C}^{2}\;:\;|zw|<1/2\}$ with the standard Euclidean ball $\mathbb{B}$ in $\mathbb{C}^{2}$. This indicates how we may construct plenty of such examples which also we shall deal later in the sequel at appropriate places; in-particular, subsequent to the next remark.
Remark 3.12.
In the setting of the foregoing proposition 1.4, being a graph, $Z$ is easily seen to be biholomorphic to $D_{L}$ as complex manifolds, in the setting of the above theorem. Thereby, it immediately follows that any pair of retracts through the origin in $D$, of the same dimension are diffeomorphic. Indeed, being star-shaped domains in $L$, they are diffeomorphic to $L$ – no pseudoconvexity is required here; the general fact being used here is that any star-shaped domain in $\mathbb{R}^{N}$ is $C^{\infty}$-diffeomorphic to $\mathbb{R}^{N}$ (exercise $7$ of chapter $8$ of [Brocker-Janich-difftop-bk]). It is therefore of interest then (as is in Several Complex Variables), to note that such a pair of retracts (through the origin of the same balanced domain $D$ and) of the same dimension despite being mutually diffeomorphic need not be biholomorphic (indeed, they are rarely so). To see this first by what is perhaps the simplest example, consider $D=\Delta^{2}\times\mathbb{B}^{2}$ as a subdomain of $\mathbb{C}^{4}$; let $Z_{1}$ denote the projection of $D$ onto the span of the first pair of standard basis vectors of $\mathbb{C}^{4}$ and $Z_{2}$, the projection of $D$ onto the span of the remaining last pair of standard basis vectors of $\mathbb{C}^{4}$. By the well-known biholomorphic inequivalence of the ball and polydisc, $Z_{1}\not\simeq Z_{2}$ as $Z_{1},Z_{2}$ are just copies of the bidisc $\Delta^{2}$ and the ball $\mathbb{B}^{2}$ respectively. To give another example, again as simple as possible but this time of a ball $D$ which is not biholomorphic to a product domain, in-fact with boundary both smooth as well as holomorphically extreme and, which contains pairs of linear retracts (thereby through the origin), of the same dimension that are biholomorphically inequivalent, consider
@@
where the exponents $p_{j}$’s are taken to be bigger than one and distinct (we disallow $N=2$ for that case is exceptional owing to the fact that all non-trivial retracts of convex domains in this case are at-most of dimension one and any pair of one-dimensional retracts are biholomorphic to $\Delta$). Then, for every $1<k<N$, there exist a pair of linear retracts of $\mathcal{E}$ of the same dimension $k$ which are diffeomorphic but biholomorphically inequivalent. This is easily seen for instance, by projecting $\mathcal{E}$ onto the span of different subsets of $k$-many of the standard basis vectors of $\mathbb{C}^{N}$ following the foregoing concrete examples.
@@
The converse of above theorem 3.9 is also true.
Proposition 3.14.
Let $D$ be a bounded balanced pseudoconvex domain and $D_{L}:=D\cap L$ be a linear retract of $D$. Let $\varphi$ be any holomorphic map from $D_{L}$ to $M$, where $M$ is a linear subspace complementary to $L$ i.e., $\mathbb{C}^{N}=L\oplus M$, is a holomorphic map with the property that $(w,\varphi(w))\in D$, whenever $w\in D_{L}$. Then ${\rm Graph}(\varphi)$, the graph of $\varphi$ over $D_{L}$ in $D$, is a retract of $D$.
Proof.
Since $D_{L}$ is a linear retract of $D$, there exists a linear projection map $P$ such that $image(P)=D_{L}$. By idempotency of $P$, we can write $\mathbb{C}^{N}=L\oplus M$, where $M=\ker(P)$ and $L={\rm image}(I-P)$. Consider the map $\psi:D\to D$ defined by $\psi(z)=(T(z),\varphi(T(z)))$; this can be verified to be idempotent by a direct computation, thereby showing that $\psi$ is a holomorphic retraction from $D$ onto ${\rm Graph}(\varphi)$ over $D_{L}$, finishing the proof. ∎
It must be noted that the above does not mean that non-linear retracts automatically come into existence by taking the graph of any holomorphic function over any given linear retract of $D$ for, the graph can fail to be contained within $D$; that this is the only obstruction to generating non-linear retracts by graphs is the content of the above proposition. However, the obstruction is substantial: for instance, even in the simple case of $\mathbb{B}$, the standard unit ball with respect to the usual $l^{2}$-norm on $\mathbb{C}^{N}$, every retract through the origin is just a linear subspace and in fact, all retracts of $\mathbb{B}$ are affine-linear, as was first established by Rudin who then exposited it in chapter $8$ of his book [Rud_book_Fn_theory_unitball] as well. The proof of this, atleast as in [Rud_book_Fn_theory_unitball], uses the special feature of the $l^{2}$-norm that it comes from an inner product; however, this fact about the linearity of retracts through the origin is actually true for every bounded balanced convex domain with $\mathbb{C}$-extremal boundary. This is of-course the result of Vesentini which we extend and give a more elementary proof as already mentioned in the introduction i.e., we now lay down the proof of theorem 1.3.
Proof of theorem 1.3: Let $\rho:D\longrightarrow D$ be a retraction and $\rho(0)=0$. By lemma 3.7, $D\rho|_{0}$ is a projection of norm $1$ from $D$ onto $D_{L}=D\cap L$, where $L=T_{0}Z$. Note that $(D\rho|_{0})_{|_{D_{L}}}:D_{L}\longrightarrow D_{L}$ is the identity map on $D_{L}$ and that $D_{L}$ is a bounded balanced convex domain in $\mathbb{C}^{N}$. It follows from the proof of Krein-Milman theorem (the reader is referred to the book [Meggn] for the standard proof we have in mind here), that $D_{L}$ must have atleast one $\mathbb{R}$-extremal boundary point, say $z_{0}$. As $D\rho(0)$ acts as the identity map on $D_{L}$ (thereby in particular $D\rho(0)(z_{0})=z_{0}$), it follows that there is atleast one $\mathbb{R}$-extremal point in $D\rho(0)(D_{L})$ namely $z_{0}$ itself! To be precise $z_{0}$ is an $\mathbb{R}$-extremal point of $D$ not just $D_{L}$. Applying theorem 3.5 to the map $\rho_{|_{D_{L}}}$ viewed as a map from $D_{L}$ into the (larger) domain $D$, finishes the proof that $Z$ is linear.
@@
$L=\Delta^{N}\cap Y=\{(F_{1}(z),\ldots,F_{N}(z)):z\in M\}.$
Observe that the $F_{j}$’s are holomorphic functions from $\Delta^{N}$ to $\Delta$ and $F_{j}(z)=G_{j}(z_{j_{1}},\ldots,z_{j_{r}})$ for some holomorphic function $G_{j}$, which is bounded by $1$. Without loss of generality assume that each $F_{i}$’s are linear, since $F(z)=(F_{1}(z),\ldots,F_{N}(z))$ restricted to $L$ is linear. So, we get
for some scalars $c_{i,j},~{}j\in J$.
@@
$w_{k}:=\begin{cases}\frac{|c_{m,j_{r}}|}{c_{m,j_{r}}},\ \text{if}\ k=j_{r}~{}\text{for some}~{}r\ \text{and}\ c_{m,j_{r}}\neq 0\\ 0,\quad\quad\text{otherwise}\end{cases}.$
@@
Hence we get $\sum_{j\in J}|c_{m,j}|\leq 1$. Recall that $M$ is spanned by $\{e_{j_{1}},\ldots,e_{j_{r}}\}$. Since $F$ is determined by the variables $z_{j_{1}},\ldots,z_{j_{r}}$, we get $Y$ is spanned by $\{F(e_{j_{1}}),\ldots,F(e_{j_{r}})\}$. By definition of $F_{i}$, we get
$F_{i}(e_{j_{k}})=\begin{cases}\delta_{i,j_{k}},\quad\qquad\text{if}~{}i\in J\\ c_{i,j_{k}},\qquad\text{if}~{}i\notin J,\\ \end{cases}$
where $\delta_{i,j}$ denotes the Kronecker delta function.
By definition of $F$, we get
$F(e_{j_{k}})=(F_{1}(e_{j_{k}}),\ldots,F_{N}(e_{j_{k}}))=e_{j_{k}}+\sum_{i\notin J}c_{i,j_{k}}e_{i}.$